5.3 is what percent of 50

Answer by rwm(914)

5.3=x*50
5.3/50=x
.106=x
10.6%=x
check reasonableness
5 /50=.10=10%
5 is 10% of 50
so it is reasonable that 5.3 is slightly more than 10%.